OpenStudy (anonymous):

Plz help Solve the perimeter formula, P=2l+2w, for w

OpenStudy (anonymous):

help

OpenStudy (godlovesme):



OpenStudy (anonymous):

?

OpenStudy (danjs):

Subtract 2*L from both sides of the equation first

OpenStudy (danjs):

P - 2L = 2L - 2L + 2W P - 2*L = 2*W

OpenStudy (godlovesme):

 or u can use danjs's method :) @zzumfelde009